Assume that Product Z is made up of four units of A and two units of C and two units of D. A is made of three units of D. The total demand for Z is 2000 units Additional information is the following:

Item Z A C D
On hand 750 500 460 2,000
Safety stock 0 120 140 200
Allocated 0 80 160 200
Scheduled Receipts -- 500 in Week 1 1,500 in Week 1 --
Provide the additional number of all items, including Z, that are required to meet the total demand shown for Product Z. Ignore the timing of these requirements
Answer #1

Requirement of Z = demand -Units on hand = 2000-750 =1250

Requirement of A = (4 x Number of Z needed) + safety stock - units on hand - units allocated - scheduled receipts

= (4x1250)+ 120 -500-80-500 = 4040

Requirement of C = ( 2x Number of Z needed) + Safety stock - units on hand - allocation -scheduled receipts

= (2x1250) +140-460-160-1500 = 520

Requirement of D = 2 x Number of Z needed + 3 x Number of A needed + safety stock - allocation -units on hand - scheduled receipts

= 2x1250 + 3x4040 +200 -200 -2000 =12620